11 11 TEX L^{A}TEX

The Belgian HOWTO

by Dag Wieërs, dag@mind.be v1.0.11, 11 September 1999

This document describes how to tweak Linux for Belgian users and lists Linux user groups, businesses and other resources in Belgium.

Contents

1 Introduction 3 1.1 Why this document? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 New versions of this document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Contributions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Feedback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Configuration 4

2.1 Keyboard setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.1.1 Console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.1.2 XFree86 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2.2 Display and applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.1 Using the ISO-8859-1 font in console . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.2 Support for the Euro symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.3 Time zone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.4 Locale support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.4.1 Dutch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4.2 French . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4.3 German . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.4.4 Walloon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.5 HOWTO translations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.5.1 Dutch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.5.2 French . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.5.3 German . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.6 Native language support and codepages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3 Setting up Internet (Belgian ISP information) 10

3.1 Dialing in with your modem (PPP) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.1.1 Configuring the modem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.1.2 Installing pppd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.1.3 pap-secrets / chap-secrets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.1.4 Network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.1.5 Dial + Hangup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.1.6 ISP-dependant information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

3.2 Dialing in with your ISDN-modem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.2.1 Updating the kernel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.2.2 ISDN4Linux toolkit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.2.3 Configuring the interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.2.4 pap-secrets / chap-secrets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.2.5 Network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.2.6 Dial + Hangup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

3.3 Cablemodem (DHCP) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.4 ADSL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4 Belgian Linux user groups 18

4.1 ALLIN Linux User Group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.2 Antwerp Linux User Group (ALUG) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.3 Brussels Linux User Group (BLUG) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.4 HCC Limburgse Linux Gebruikers Groep (HLLGG) . . . . . . . . . . . . . . . . . . . . . . . 20 4.5 HCC Linux Gebruikers Groep Antwerpen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.6 HCC Linux Gebruikers Groep Leuven . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.7 Independent Group of Unix-Alikes and Network Activists (IGUANA) . . . . . . . . . . . . . 21 4.8 Infogroep (IGWE) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.9 Leuvense Linux Users (L2U) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.10 Limburgse Linux User Group (LILUG) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.11 Linux Belgian Team . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.12 Linux Gebruikers Oost-Vlaanderen (LGOVL) . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.13 Linux User Group De Bruxelles (BxLUG) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.14 Linux User Group De Namur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.15 Linux User Group Mons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.16 Linux User Movement Underground Mad Belgian Aliens (LUMUMBA) . . . . . . . . . . . . 24 4.17 linuxbe.org . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.18 Local Linux Support Group Gent (LLSGG) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.19 Open Technology Assembly Linux Special Interest Group . . . . . . . . . . . . . . . . . . . . 26 4.20 Student Information Networking (SIN) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.21 TINA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.22 Unix Lovers Yield Student Services & Internet Support (ULYSSIS) . . . . . . . . . . . . . . . 27 4.23 Zeus WPI - Werkgroep Informatica RUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

5 Belgian Linux businesses 28 5.1 aDOC Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.2 Better Access N.V. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.3 D. Connect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.4 Double Barrel Consultancy and Productions . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.5 grmbl productions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 5.6 LIFE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 5.7 Phidani Software SPRL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 5.8 Spier BVBA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

5.9 VirgoPlus sprl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

6 Belgian Linux resources 32

6.1 Websites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 6.1.1 Linux in Belgium . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 6.1.2 Dutch Linux-sites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 6.1.3 French Linux-sites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 6.1.4 Mirrors in Belgium . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

6.2 HOWTO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.2.1 Web . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.2.2 FTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

6.3 FTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.3.1 Kernel-sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.3.2 Distributions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.3.3 Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

6.4 Usenet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 6.5 Mailinglists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

6.5.1 Linux user groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 6.6 IRC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

1 Introduction

1.1 Why this document?

Since there was no information anywhere on the net for Belgian linux users, I started to collect stuff that I thought could be important. This HOWTO aims to be the definitive list of all Belgian-related Linux information. If you think not, you have no excuse to complain and certainly not to not contribute. ;)

1.2 New versions of this document

If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the supervisor Tim Bynum linux-howto@metalab.unc.edu. Tim Bynum will post the listing to several national and international newsgroups on a monthly basis. In addition, the Belgian HOWTO can be found on the World Wide Web at <http://dag.mind.be/howto/> New versions of the Belgian HOWTO are always placed at this site first, so please be sure to check if the copy you are reading is still up to date!

1.3 Contributions

A few people mailed me their suggestions and improvements, thanks go to:

• Wim Vandeputte wvdeputte@reptile.rug.ac.be

• Pablo Saratxaga srtxg@ping.be

• Christophe Lambin clambin@skynet.be

• Geert Uytterhoeven geert@cs.kuleuven.ac.be

Many thanks go to Ivo Clarysse soggie@iguana.be who started something similar earlier () and his project helped me a lot to get this document online.

1.4 Feedback

If some information seems to be wrong, deceptive or missing, I’d appreciate if you mailed me the improvements. Since I’m just human this document isn’t bug-free, but your contribution can and will make a difference.

To add yourself to the Linux user groups or businesses, collect all necessary information and mail it to me.

2 Configuration

2.1 Keyboard setup

Linux provides 2 ways to set up your keyboard. At the console you can use loadkeys and under XFree86 you can use xmodmap.

2.1.1 Console

To use the keytable for a Belgian keyboard you can use

        loadkeys be2-latin1

or

        loadkeys be-latin1

The only difference between the two is that be2-latin1 add support for twosuperior and threesuperior (keycode 41) and grave (keycode 43).

Usually loadkeys is started at boottime from the scripts located somewhere in /etc/rc.d Different distributions handle it differentely. But most distributions have a simple program (like kbdconfig in RedHat or yast in SuSE) to change the scripts.

2.1.2 XFree86

To set your keyboard properly under XFree86 you’ve got more choices. By running Xconfigurator (or a similar program) a config-file for XFree86 is created (with some standard content). This file is called XF86Config and is usually located in /etc or /etc/X11

Make sure (and change otherwise) that something similar to this is given under the Keyboard Section.

        Section "Keyboard"
                Protocol        "Standard"
                XkbRules        "xfree86"
                XkbModel        "pc101"
                XkbLayout       "be"
        EndSection

If you own a ’microsoft’ keyboard (or the less-known penguin keyboards ;p) you can enable those extra 3 keys by changing the "pc101" into "pc104". (I use the extra keys as meta-keys in my windowmanager so some shortcuts doesn’t conflict with the internal ones of my wm).

To get the AltGr key to work under XFree86 simply add in the Keyboard Section:

                RightAlt        ModeShift

To get some key-combinations working under X11, you might want to change in file /usr/X11R6/lib/X11/xkb/symbols/be:

        key <AD11> {    [ dead_circumflex,       diaeresis      ],
                        [     bracketleft                       ]       };
        key <AC11> {    [          ugrave,         percent      ],
                        [      quoteright                       ]       };
        key <BKSL> {    [              mu,        sterling      ],
                        [       quoteleft                       ]       };
        key <AB10> {    [           equal,            plus      ],
                        [      asciitilde                       ]       };

by

        key <AD11> {    [ dead_circumflex,  dead_diaeresis      ],
                        [     bracketleft                       ]       };
        key <AC11> {    [          ugrave,         percent      ],
                        [      dead_acute                       ]       };
        key <BKSL> {    [              mu,        sterling      ],
                        [      dead_grave                       ]       };
        key <AB10> {    [           equal,            plus      ],
                        [      dead_tilde                       ]       };

Some people prefer to use a ’compose’-key to enter their special characters, like:

<compose>'e => e
<compose>/o => oe
<compose>c, => c
<compose>co => (C)
<compose>ss => ss

Solution:

xmodmap -e "keycode xx = Multi_key"

or

xmodmap -e "keysym yy = Multi_key"

and ’xx’ is the keycode and ’yy’ the kesym of the key you choose to be the ’compose’-key. (use xev to get these values)

Alternatively you can add

keycode xx = Multi_key

to your /.xmodmaprc. This way it happens automagically.

2.2 Display and applications

Some applications have to be compiled as 8-bit clean to work well with the European characterset. Others like to be told in advance.

2.2.1 Using the ISO-8859-1 font in console

You can load the ISO-8859-1 font by typing the following in console:

setfont lat1u-16.psf
mapscrn trivial
echo -ne '\033(K'

To be able to use ë, è, é or ç in console, you might want to add these to your /.inputrc:

set meta-flag on
set convert-meta off
set output-meta on

(this must be done for every user, if you want this to be default for all users you can either add this file to /etc/skel/ or you could add these lines to /etc/inputrc and add the following line to /etc/bashrc,

        export INPUTRC=/etc/inputrc

)

Note: this is only useful to programs that use readline (like bash) Also set the following environment variables:

        LC_CTYPE=iso-8859-1
        SYSFONT=lat1u-16.psf
        SYSTERM=linux

2.2.2 Support for the Euro symbol

Since Europe had to invent a new character, the Euro, there’s a new standard. It’s called ISO-8859-15 (aka Latin-9 or Latin-0). To add support you need both a new font and new keymaps. They’re included with the EURO-package and can be found at: <ftp://ftp.freshmeat.net/pub/euro-patch/>. Soon this will be part of every major distribution.

The changes in the keymap-file consist of the following lines:

        keycode  18 = +e        +E      currency
        keycode  46 = +c        +C      cent

And the following environment-variable should be set:

        SYSFONT=lat0-16.psf

to make sure you’re using the right font.

2.3 Time zone

Since Belgium is located in the Central European Time zone (aka MET) which (in the winter) is equivalent to the Greenwhich Mean Time plus 1 (GMT+1), you can simply link /usr/lib/zoneinfo/localtime/MET to /etc/localtime symbolicly like:

        ln -sf /usr/lib/zoneinfo/MET /etc/localtime

or

        ln -sf /usr/lib/zoneinfo/Europe/Brussels /etc/localtime

This automagically sets Daylight Savings (which is GMT+2 in the summer).

Note: that different distributions have different paths, zoneinfo can be located in /usr/share or such...

Note2: RedHat uses a tool called timeconfig, SuSE uses yast

To change the CMOS clock and then synchronize the system time with the CMOS clock, do something like this:

        hwclock --set --date="Feb 25 03:38"
        clock -u -s

If your clock is set to local time (which is discouraged but if you have other OS’s installed a must), you can do:

        hwclock --set --date="Feb 25 04:38"
        clock -s

[Noticed the 1 hour difference ? ;p]

2.4 Locale support

Not much programs support locale yet, if you want to know more about locale or how to write your own programs to support it, check: [put the damn url here ;p]

Be aware that locale makes it harder to resolve problems as less people can help you !! You might understand a problem when it’s in technical English, but what if it’s translated in some obscure French ?!? Some obviously won’t understand both. ;p

2.4.1 Dutch

To enable support for the Dutch locale in Belgium on a system with locale support you just have to set the following environment variables:

        LANG=nl
        LANGUAGE=nl_BE

Try nl_BE.ISO_8859-1 if nl_BE does not work.

2.4.2 French

If you want French locale support in Belgium, set the following environment variabels:

        LANG=fr
        LANGUAGE=fr_BE

Try fr_BE.ISO_8859-1 if fr_BE does not work.

2.4.3 German

If you want instead German locale support in Belgium, set the following environment variabels:

        LANG=de
        LANGUAGE=de_BE

Try de_BE.ISO_8859-1 if de_BE does not work.

2.4.4 Walloon

For Walloon locale support you can find files and information at: <http://www.ping.be/linux/walon/>. To use it set:

        LANG=wa
        LANGUAGE=wa_BE:fr_BE

Try wa_BE.ISO_8859-1 if wa_BE does not work.

2.5 HOWTO translations

As Dutch is less wide-spread, it is the least supported. But work is being done to resolve this ;) (and maybe *you* can help ?!?) French and German translations obviously are already taken care of.

2.5.1 Dutch

A succesfull initiative is making progress, but there’s still a lot to do. If you want to help, start translating

and send it there. Check: <http://nl.linux.org/doc/HOWTO/>.

2.5.2 French

French translations of HOWTO’s can be found at: <ftp://metalab.unc.edu/pub/Linux/docs/HOWTO/
translations/french/>

2.5.3 German

German translations of HOWTO’s can be found at: <ftp://metalab.unc.edu/pub/Linux/docs/HOWTO/
translations/german/>

2.6 Native language support and codepages

These options add support for the Belgian characterset and codepages. If you’d like to have these, compile a kernel with these options put on. They only apply to some filesystems (FAT-based, ISO9660, ...).

        < > Codepage 850 (Europe)
        < > NLS ISO 8859-1  (Latin 1; Western European Languages)
        < > NLS ISO 8859-15 (Latin 9; Western European Languages with Euro)

Located in Filesystems > Native Language Support

If your kernel is a modular kernel that came with your distribution (=you never compiled a kernel before), then you might want to type:

        insmod nls_cp850
        insmod nls_iso8859_1
        insmod nls_iso8859_15

3 Setting up Internet (Belgian ISP information)

3.1 Dialing in with your modem (PPP)

To configure PPP correctly, you should check the PPP-HOWTO <http://www.interweft.com.au/other/ ppp-howto/ppp-howto.html>. I try to explain it briefly, but if you have any questions or problems, read the HOWTO very carefully.

3.1.1 Configuring the modem

To configure your modem, you can alter the initstring of your modem. To learn more about your modem, check Ask Mr. Modem <http://www.spy.net/~dustin/modem/>. Beware, if you have a winmodem you will not get it to work. More information about winmodems is at: <http://linmodems.org/>

3.1.2 Installing pppd

Now, you need to have the ppp-package installed (check this by typing pppd –version), make sure you have a recent one. Then you have to make sure you’ve got these files:

/usr/local/bin/ppp

        #!/bin/sh
        case $1 in
                on|start)
                TELEPHONE='555-1212'    # The telephone number for the connection
                ACCOUNT='your_login'    # The account name for logon (as in 'George Burns')
                PASSWORD='your_pass'    # The password for this account (and 'Gracie Allen')
                INITSTRING='AT'         # Modem initstring
                IPLOCAL=0.0.0.0         # Local IP address if known. Dynamic = 0.0.0.0
                IPREMOTE=0.0.0.0        # Remote IP address if desired. Normally 0.0.0.0
                DEVICE=/dev/ttyS1       # Serial Device com1=ttyS0, com2=ttyS1,...
                SPEED=57600             # 19200, 38400 or 57600 (don't try something different)

export TELEPHONE ACCOUNT PASSWORD INITSTRING
exec /usr/sbin/pppd $DEVICE $SPEED $IPLOCAL:$IPREMOTE \
        user $ACCOUNT connect /etc/ppp/ppp-on-dialer
;;
off|stop)
if [ -r /var/run/ppp0.pid ]; then
        kill -INT `cat /var/run/ppp0.pid`
        if [ ! "$?" = "0" ]; then
                rm -f /var/run/ppp0.pid
                echo "ERROR: Removed stale pid file"
                exit 1
        fi
        exit 0
fi
echo "ERROR: PPP link is not active on ppp0"
exit 1
;;
info|status)
/usr/sbin/pppstats
;;
*)
echo "Usage: ppp on|off|info|start|stop|status";
;;
esac

/etc/ppp/ppp-on-dialer

#!/bin/sh
exec /usr/sbin/chat -v                                  \
TIMEOUT         3                               \
ABORT           '\nBUSY\r'                      \
ABORT           '\nNO ANSWER\r'                 \
ABORT           '\nNO CARRIER\r'                \
ABORT           '\nNO DIALTONE\r'               \
ABORT           '\nRINGING\r\n\r\nRINGING\r'    \
''              "\r$INITSTRING\r"               \
'OK-+++\c-OK'   'ATH0'                          \
TIMEOUT         30                              \
OK              "ATDT$TELEPHONE"                \
CONNECT         ''                              \
ogin:--ogin:    "$ACCOUNT"                      \
ord:            "$PASSWORD"

You can remove the -v once everything looks normal.

The pppd command uses /etc/ppp/options to list its options, change these options to whatever fits. These defaults normally work in most cases.

asyncmap 20a0000
crtscts
debug
default-mru
defaultroute
detach
escape 11,13,ff
hide-password
ipcp-accept-local
ipcp-accept-remote
lcp-echo-failure 4
lcp-echo-interval 400
modem
mtu 1500
netmask 255.255.255.0
noipdefault
passive
#idle 300
#kdebug 0
#lock
#-vj

Once dialing in works like a charm, you can leave debug out.

3.1.3 pap-secrets / chap-secrets

If your ISP requires CHAP or PAP, you need one of these files with their proper username and password:

/etc/ppp/pap-secrets or /etc/ppp/chap-secrets

#client         server  secret          IP
your_login      *       your_pass       *

3.1.4 Network

To configure your nameservers, change /etc/resolv.conf . If dialing in is succesfull, but you can’t get it to connect to URL’s: this is probably the cause.

domain domain_of_your_isp
nameserver your_primary_nameserver
nameserver your_secondary_nameserver

3.1.5 Dial + Hangup

Then if /usr/local/bin is in your PATH, you can easily dial in with:

        ppp on

and hangup with:

        ppp off

and to get some statistics, you do:

        ppp status

3.1.6 ISP-dependant information

Digibel:

Web:

         <http://www.digibel.org/>

Eunet:

Authentication: text-based

Options:

proxyarp Web:

         <http://www.eunet.be/>

Euronet (Wannadoo):

Authentication: text-based

Web:

         <http://www.euronet.be/>

Online:

Authentication: text-based

Options:

noccp Web:

    <http://www.online.be/>

Ping:

Authentication:

text-based Web:

    <http://www.ping.be/>

Planet Internet:

Authentication:

CHAP Web:

    <http://www.planetinternet.be/>

Skynet (Belgacom):

Authentication:

PAP Options:

defaultroute Web:

    <http://www.skynet.be/>

UUNet:

Authentication:

PAP Web:

    <http://www.uunet.be/>

XS4ALL:

Authentication:

PAP Web:

    <http://www.xs4all.be/>

3.2 Dialing in with your ISDN-modem

Contributed by Christophe Lambin clambin@skynet.be This section does not aim to be a definitive guide on ISDN for Linux. For a more detailed discussion on the topic, see Paul Slootman’s ISDN4Linux HOWTO and ISDN4Linux-FAQ.

3.2.1 Updating the kernel

If you’re using a 2.2 kernel, it’s recommended to get an updated version of the ISDN code. You can

retrieve the CVS at <ftp://ftp.suse.com/pub/isdn4linux/>. For more information, check: <http:
//www.brisse.dk/site/linux/docs/isdn.htm>

Warning: if you’ve recently purchased an Eicon Diva 2.0 PCI (eg, via Belgacom), there’s a good chance you’ve actually got a Diva 2.01.In this case, you must get a version dated 1 July 1999 or later !

Once the ISDN is in place, you can start up the driver. Eg:

modprobe hisax type=11 protocol=2

Refer to the Readme’s for the right parameters and values for your card.

3.2.2 ISDN4Linux toolkit

Next, you’ll need the ISDN4Linux toolkit. You can retrieve the latest version of the toolkit at: <ftp:

//ftp.suse.com/pub/isdn4linux/>

3.2.3 Configuring the interface

With the tools installed and configured, write a script to configure the interface, to be used for the ISDN connection. As always in Linux, there’s no one correct way of doing this. I’ve put these in a script

/etc/rc.d/rc.isdn:

#!/bin/sh
MSNREMOTE='555-1212'    # Phone number of ISP
MSNLOCAL='555-1313'     # my number, without 0, with areacode
ACCOUNT='george'        # The account for logon (as 'George Burns')
IPLOCAL=10.0.0.2        # my fixed IP (use 10.0.0.2 if no fixed)
IPREMOTE=0.0.0.0        # IP number of ISP
INTERFACE=ippp0
/sbin/modprobe hisax type=11 protocol=2
/sbin/isdnctrl verbose 3 system on
/sbin/isdnctrl addif $IF
/sbin/isdnctrl secure $IF on
/sbin/isdnctrl addphone $IF out $MSNREMOTE
/sbin/isdnctrl eaz $IF $MSNLOCAL
/sbin/isdnctrl huptimeout $IF 300
/sbin/isdnctrl l2_prot $IF hdlc
/sbin/isdnctrl l3_prot $IF trans
/sbin/isdnctrl encap $IF syncppp
#       /sbin/isdnctrl dialmode $IF auto
/sbin/ifconfig $IF $IPLOCAL pointopoint $IPREMOTE -arp -broadcast
/sbin/ipppd /dev/ippp0 user $ACCOUNT $IPLOCAL:$IPREMOTE

To start this at boot time, make it executable and append the following to /etc/rc.d/rc.local:

if [ -x /etc/rc.d/rc.isdn ]; then
        . /etc/rc.d/rc.isdn
fi

The ipppd command gets its parameters passed through a file, /etc/ppp/ioptions:

-ac
-bsdcomp
debug
defaultroute
ipcp-accept-local
ipcp-accept-remote
mru 1524
mtu 1500
noipdefault
-pc
useifip
-vj
-vjccomp
#idle 360
#persistent

Warning: do NOT specify +pap or +chap in this file. This specifies the authentication that ipppd should use for an INCOMING client. If you were to use this to connect to your ISP, ipppd would wait for the ISP to authenticate itself using the specified protocol.

Finally, create /etc/ppp/ip-down.local to handle the shutdown of the interface:

#!/bin/sh
/sbin/ifconfig $1 down
sleep 1
/sbin/ifconfig $1 10.0.0.2 pointopoint

Note: this is to handle some problems with routes on shutdowns. Anyone know of a clean(er) solution ?

3.2.4 pap-secrets / chap-secrets

If your ISP uses PAP or CHAP, create a file /etc/ppp/pap-secrets or /etc/ppp/chap-secrets. Its format is:

#client         server  secret          IP
your_login      *       your_pass       *

3.2.5 Network

Not really related to ISDN, but a lot of people forget this step (too eager to try out all the stuff they typed in above ? :-)).

Anyway, you need to configure the DNS, by creating a file /etc/resolv.conf :

        domain your_domain_of_your_isp
        nameserver your_primary_nameserver
        nameserver your_secondary_nameserver

3.2.6 Dial + Hangup

Finally, you can write a little wrapper to start and stop the ISDN connection. I’ve put this as /usr/local/bin/isdn:

        #!/bin/bash
        case $1 in
                on|start)
                /sbin/isdnctrl dial ippp0
                ;;
                off|stop)
                /sbin/isdnctrl hangup ippp0
                ;;
                info|status)
                /sbin/isdnctrl list ippp0
                ;;
                *)
                echo "Usage: isdn on|off|info|start|stop|status"
                ;;
        esac

Then if /usr/local/bin is in your PATH, you can easily dial in with:

        isdn on

and hangup with:

        isdn off

3.3 Cablemodem (DHCP)

DHCP was designed to make life easier, and most of the times it does ;) More information to set up your

DHCP client can be found at: <http://metalab.unc.edu/LDP/HOWTO/mini/DHCP.html> which describes

everything you need to know. (along with the information below)

Lots of people want to connect a whole network to their cablemodem, that’s not a problem. You’ll need IP-

Masquerading ( <http://metalab.unc.edu/LDP/HOWTO/mini/IP-Masquerade.html>) and make sure the

interface (that speaks to the cablemodem) has the right MAC-address !! Another solution (in case of problems) is to reset your cablemodem.

Chello (TVD):

Web:

         <http://www.chello.be/>

More information:

         <http://foobar.starlab.net/~soggie/tvd_linux/>

Pandora (Telenet):

Automatic proxy:

http://pac.pandora.be:8080 http proxy:

export http_proxy="http://proxy.pandora.be:8080/" Web:

         <http://www.pandora.be/>

3.4 ADSL

Turboline (Belgacom):

Some things:

using DANA Web:

         <http://www.turboline.be/>

More information:

         <http://foobar.starlab.net/~soggie/turboline/>

4 Belgian Linux user groups

More information can be found in the User Group HOWTO located at: <http://www.ntlug.org/

~kclark/> A complete list of Linux User Groups is at: <http://www.ssc.com/linux/glue/> and at:
<http://limestone.uoregon.edu/woven/lugww/>

4.1 ALLIN Linux User Group

        Mechelbaan 260
        B-3130 Begijnendijk

Email:

allin@linuxfreak.com Web:

     <http://come.to/allin/>

Contact:

Paul Terweduwe paul@faes.be Meetingplace:

’T Paviljoen in Heist Op Den Berg Last modified:

13 April 1999.

4.2 Antwerp Linux User Group (ALUG)

        Van Luppenstraat 70
        B-2018 Antwerp

Phone:

+32 3 218 63 54 Email:

fv.alug@advalvas.be Web:

     <http://sunsite.belnet.be/alug/>

Contact:

Patrick Op de Beeck poe@glo.be Meetings:

Almost every friday at 19h, see homepage first ! Meetingplace:

Hof van Rieth, Molenlei 68, Mortsel / Cultureel Centrum Berchem, Driekoningenstraat 126 Last modified:

8 March 1999.

4.3 Brussels Linux User Group (BLUG)

        Brussels Linux User Group (BLUG)

Email:

johanb@linux.rave.org Web:

     <http://linux.rave.org/>

Last modified:

24 February 1999.

4.4 HCC Limburgse Linux Gebruikers Groep (HLLGG)

        HCC Limburgse Linux Gebruikers Groep

Email:

Linuxgg@hotmail.com Web:

     <http://www.uunet.be/hcc/Limburg/Linuxgg>

Contact:

Andy Peeters & Bruno Tony Last modified:

31 May 1999.

4.5 HCC Linux Gebruikers Groep Antwerpen

        Armand Verachtert
        Turnhoutsebaan 421
        2140 Borgerhout

Web:

     <http://users.glo.be/~poe/index.html>

Last modified:

25 February 1999.

4.6 HCC Linux Gebruikers Groep Leuven

        Diegemstraat 61
        B-1930 Zaventem

Email:

Yvo.Dries@village.uunet.be Web:

     <http://www.uunet.be/hcc/>

Contact:

Yvo Dries Yvo.Dries@village.uunet.be Last modified:

2 August 1999.

4.7 Independent Group of Unix-Alikes and Network Activists (IGUANA)

        Milcampslaan 101
        B-1040 Schaarbeek

Phone:

+32 2 657 77 36 Fax:

+32 2 657 77 38 Email:

info@iguana.be Web:

     <http://www.iguana.be/>

Last modified:

5 March 1999.

4.8 Infogroep (IGWE)

        Pleinlaan 2, 5F218 (building F)
        B-1050 Brussels

Fax:

+32 2 629 33 89 Email:

ig@igwe.vub.ac.be Web:

     <http://igwe.vub.ac.be/>

Last modified:

6 July 1999.

4.9 Leuvense Linux Users (L2U)

        Leuvense Linux Users (L2U)

Email:

darkcompanion@village.uunet.be Web:

     <http://gallery.uunet.be/darkcompanion/linux/>

Last modified:

23 February 1999.

4.10 Limburgse Linux User Group (LILUG)

        Limburgse Linux User Group (LILUG)

Email:

frank.eersels@ping.be Web:

     <http://lilug.linuxatwork.at/>

Contact:

Frank Eersels frank.eersels@ping.be Last modified:

6 May 1999.

4.11 Linux Belgian Team

        Rue d'Aineffe 4
        B-4317 Borlez Faimes

Email:

manu@rtfm.be Web:

     <http://linux.rtfm.be/>

Contact:

Alexandre J.D. Dulaunoy adulau@rtfm.be Last modified:

25 February 1999.

4.12 Linux Gebruikers Oost-Vlaanderen (LGOVL)

        Linux Gebruikers Oost-Vlaanderen

Email:

dennis@lgovl.freeservers.com Web:

     <http://lgovl.freeservers.com/>

Last modified:

25 February 1999.

4.13 Linux User Group De Bruxelles (BxLUG)

        Linux User Group De Bruxelles (BxLUG)

Email:

bxl_information@linuxbe.org Web:

     <http://bxlug.linuxbe.org/>

Contact:

Raphael Bauduin bxl_information@linuxbe.org Last modified:

17 July 1999.

4.14 Linux User Group De Namur

        Linux User Group De Namur

Email:

namurlugs@linuxbe.org Web:

     <http://namur.linuxbe.org/>

Last modified:

17 July 1999.

4.15 Linux User Group Mons

        Linux User Group Mons

Email:

webmaster@linuxwallonie.org Web:

     <http://www.linuxwallonie.org/>

Last modified:

26 May 1999.

4.16 Linux User Movement Underground Mad Belgian Aliens (LUMUMBA)

        Universiteitslaan 1, Filii Lamberti
        3590 Diepenbeek

Email:

helpdesk@lumumba.luc.ac.be Web:

     <http://lumumba.luc.ac.be/>

Last modified:

25 February 1999.

4.17 linuxbe.org

        linuxbe.org

Email:

info@lists.linuxbe.org Web:

     <http://linuxbe.org/>

Contact:

Cedric Gavage niddle@linuxbe.org Last modified:

5 March 1999.

4.18 Local Linux Support Group Gent (LLSGG)

        Wim Vandeputte
        PB 144
        B-9000 Gent 12

Phone:

+32 9 240 29 17 Fax:

+32 9 240 34 39 Email:

wvdputte@reptile.rug.ac.be Web:

     <http://llsgg.rug.ac.be/>

Contact:

Wim Vandeputte wvdputte@reptile.rug.ac.be Last modified:

4 March 1999.

4.19 Open Technology Assembly Linux Special Interest Group

        Kruipstraat 14
        B-1850 Grimbergen

Phone:

+32 11 27 58 98 Fax:

+32 11 27 03 89 Email:

info@ota.be Web:

     <http://www.ota.be/sigs/linux/>

Contact:

Jan Vanhercke jan.vanhercke@c-cure.be Last modified:

4 March 1999.

4.20 Student Information Networking (SIN)

        Kleinhoefstraat 4
        B-2440 Geel

Email:

sin@sin.khk.be Web:

     <http://www.sin.khk.be/>

Last modified:

12 June 1999.

4.21 TINA

Email:

vraagtina@digibel.org Web:

     <http://tina.digibel.org/>

Contact:

Roman Pauwels art@ping.be Last modified:

3 September 1999.

4.22 Unix Lovers Yield Student Services & Internet Support (ULYSSIS)

        Blijde Inkomsstraat 21, lokaal 01.17
        B-3000 Leuven

Phone:

+32 16 32 48 61 Email:

ulyssis@ulyssis.student.kuleuven.ac.be Web:

     <http://www.ulyssis.student.kuleuven.ac.be/org/ulyssis/>

Last modified:

25 February 1999.

4.23 Zeus WPI - Werkgroep Informatica RUG

        Krijgslaan 281 S9
        B-9000 Gent

Phone:

+32 09 264 47 51 Email:

info@zeus.rug.ac.be Web:

     <http://www.zeus.rug.ac.be/>

Comment:

Restricted to RUG students Last modified:

22 June 1999.

5 Belgian Linux businesses

The Linux Documentation Project contains a worldwide list of Linux consultants, Linux Consultants HOW-TO, and can be found at: http://consult.cyrius.com/ <http://consult.cyrius.com/>.

5.1 aDOC Services

        Avenue du Pesage, 31-33
        B-1050 Brussels

Phone:

+32 2 646 00 76 Email:

sales@adoc-services.com Web:

     <http://www.adoc-services.com/>

Last modified:

25 February 1999.

5.2 Better Access N.V.

        Geldenaakse Vest 6
        3000 Leuven

Phone:

+32 16 29 80 45 Fax:

+32 16 29 80 46 Email:

info@ba.be Web:

     <http://www.ba.be/>

Type of support:

After an onsite installation, we mainly support our customers with SSH remote administration. Most problems are reported by E-mail or by the inhouse-developped webbased supportsystem. This doesn’t mean that once in a while a supportcall comes in.

Special expertise:

Security, systemadministration and setups, security (firewalling, tigerteaming, VPN, etc.) Networkdesign, wireless networking, Troubleshooting, product development, etc. Please visit http://www.ba.be for more info.

Last modified:

7 March 1999.

5.3 D. Connect

   Av. Van Goidtsnoven 33
   B-1180 Uccle

Email:

webmaster@dconnect.be Web:

<http://www.dconnect.be/>

Last modified:

25 February 1999.

5.4 Double Barrel Consultancy and Productions

   Sportstraat 28
   B-9000 Gent

Phone:

+32 9 222 77 64 Fax:

+32 9 222 49 76 Email:

mvergall@double-barrel.be Web:

<http://www.double-barrel.be/>

Contact:

Michael C. Vergallen mvergall@double-barrel.be Type of support:

phone, email remote network admin, inhouse, onsite.

Special expertise:

installation, system admin. internet and intranet connectivity, sendmail, firewalls, proxys, nameservers, troubleshooting software & hardware.

Last modified:

5 March 1999.

5.5 grmbl productions

   Korte Vuldersstraat
   B-8000 Brugge

Phone:

+32 50 344366 Email:

info@grmbl.com Web:

<http://www.grmbl.com/>

Last modified:

2 March 1999.

5.6 LIFE

   Mgr. Vanwaeyenberglaan 6/10
   B-3000 Leuven

Phone:

+32 477 57 50 71 Email:

info@life.be Web:

<http://www.life.be/>

Contact:

Jasper Nuyens jasper@life.be Type of support:

LIFE is the first company in Europe to provide 24/7 professional support for Linux in Europe. LIFE also provides server and desktop installation services, educational services and fixed price development. Exclusive on the Linux operating system.

Special expertise:

Kernel driver development, database interconnectivity (VPN, wireless networking), security auditing (firewall, VPN, ethical hacking), e-commerce. See our website for more.

Last modified:

20 April 1999.

5.7 Phidani Software SPRL

        Rue de l'autonomie 1
        B-1070 Brussels

Email:

info@phidani.be Web:

     <http://www.phidani.be/>

Last modified:

25 February 1999.

5.8 Spier BVBA

        Knaptandstraat 96-98
        B-9100  Sint-Niklaas

Email:

info@spier,be Web:

     <http://www.spier.be/>

Last modified:

02 March 1999.

5.9 VirgoPlus sprl

        Rue Dartois 1
        B-4000 Liege

Fax:

+32 4 253 00 49 Email:

contact@virgoplus.com Web:

     <http://www.virgoplus.com/>

Contact:

Bruno Mairlot bruno@virgoplus.com Type of support:

By phone, email, or remote networked administration, inhouse, onsite Special expertise:

Installation, configuration, administration of Linux, internet connection, intranet developpment, firewalls, samba, netatalk, firewalls, proxys, DNS, web server (Apache), mod_perl, MySQL, Gui interface developpment (GTK+). We have acquired a special expertise with the Red Hat distributions.

Last modified:

28 June 1999.

6 Belgian Linux resources

6.1 Websites

6.1.1 Linux in Belgium

   • Linux in Belgium <http://linbel.iguana.be/>
   • Belgian HOWTO <http://dag.mind.be/howto/Belgian-HOWTO.html>

6.1.2 Dutch Linux-sites

   • <http://nl.linux.org/doc/>

• Nederlandstalige Linux Documentatie Project <http://doc.nllgg.nl/products/>

   • Linux Support Website <http://www.linux-nl.cx/>
   • Beginnen met Linux <http://user.online.be/~koendb/linux_beginners.html>
   • Installatie Handleiding RedHat 5.2 <http://user.online.be/~koendb/redhat.html>

6.1.3 French Linux-sites

   • linuxbe.org <http://linuxbe.org/>

6.1.4 Mirrors in Belgium

• Freshmeat at Chello (TVD) <http://be.eu.mirrors.freshmeat.net/>
• Linuxberg at Wallonie Internet <http://winbe.linuxberg.com/>
• Linuxberg at Online <http://online.linuxberg.com/>
• Linuxberg at Pandora (Telenet) <http://linuxberg.pandora.be/>
• Linux Gazette at Belnet <http://sunsite.belnet.be/alug/ftp_mirrors/linux_gazette/>

6.2 HOWTO

The most important collection of Linux information on the net. Please if you encounter erroneous information in one of them, do contact the author. Only that way it will be corrected and we all stand together ;)

6.2.1 Web

• HOWTO’s at KULeuven <http://www.kulnet.kuleuven.ac.be/LDP/HOWTO/>
• HOWTO’s at SIN <http://www.sin.khk.be/mirror/LDP/>

6.2.2 FTP

• HOWTO’s at Belnet <ftp://ftp.belnet.be/mirror/sunsite.unc.edu/pub/Linux/docs/HOWTO/
  INDEX.html>

6.3 FTP

Most of the times, it doesn’t matter where you got it from, as long as it works. But to save you some time and save us some bandwidth, it’s much better to get it near you ;) Naturally I forgot some, please send your additions or corrections.

6.3.1 Kernel-sources • Official Mirror kernel-sources at Chello (TVD) <ftp://ftp.be.kernel.org/>

• Mirror kernel-sources at linux.be.eu.org <ftp://ftp.linux.be.eu.org/mirrors/ftp.kernel.org/>
• Mirror kernel-sources at Digibel <ftp://ftp.digibel.be/pub/mirrors/ftp.kernel.org/>

6.3.2 Distributions

• Red Hat

    – Official Mirror at Belnet <ftp://ftp.belnet.be/mirror/ftp.redhat.com/>
    – Mirror at Chello (TVD) <ftp://ftp.chello.be/mirror/ftp.redhat.com/>
    – Mirror at linux.be.eu.org <ftp://ftp.linux.be.eu.org/mirrors/ftp.redhat.com/>

– Mirror at Pandora (Telenet) <ftp://mirrors.pandora.be/mirrors/ftp.redhat.com/>
– Mirror at Digibel <ftp://ftp.digibel.be/pub/mirrors/ftp.redhat.com/>

• Debian

– Official Mirror at KULeuven <ftp://ftp.kulnet.kuleuven.ac.be/pub/mirror/ftp.debian.
  org/>
– Official Mirror at spnet <ftp://debian.spnet.net/>
– Mirror at EUnet <ftp://ftp.eunet.be/pub/linux/debian/>
– Mirror at Skynet <ftp://ftp.skynet.be/mirror/linux/distributions/debian/>
– Mirror at Pandora (Telenet) <ftp://mirrors.pandora.be/mirrors/ftp.debian.org/>
– Mirror at Interpac <ftp://ftp.interpac.be/mirror/linux/distributions/debian/>

• Slackware

– Official Mirror at Interpac <ftp://ftp.interpac.be/mirror/linux/distributions/
  slackware/>
– Mirror at Skynet <ftp://ftp.skynet.be/mirror/linux/distributions/slackware/>
– Mirror at Globe <ftp://ftp.glo.be/pub/mirror/Slackware-3.6/>

• SuSE

– Mirror at Pandora (Telenet) <ftp://mirrors.pandora.be/mirrors/>

• Mandrake

– Mirror at Belnet <ftp://ftp.belnet.be/packages/mandrake/>

6.3.3 Software

• Mirrors at Belnet <ftp://ftp.belnet.be/mirror/>
• Mirrors at Chello (TVD) <ftp://ftp.chello.be/mirror/>
• Mirrors at linux.be.eu.org <ftp://ftp.linux.be.eu.org/mirrors/>
• Mirrors at Pandora (Telenet) <ftp://mirrors.pandora.be/mirrors/>
• Mirrors at Digibel <ftp://ftp.digibel.be/pub/mirrors/>

6.4 Usenet

• <news:be.comp.os.linux> at WebWatch <http://www.webwatch.be/forums/be.comp.os.linux/
MainTable.html>

6.5 Mailinglists

6.5.1 Linux user groups • BLUG Mailinglist

• L2U Mailinglist

• LILUG Mailinglist

• LGOVL Mailinglist

6.6 IRC

The fastest medium to get support is IRC, but beware. Because IRC is rather anonymous it’s also easy to fool people. On IRCnet there’s a channel called #linux.be that is closely related to bcol and the Belgian Linux-scene. The channel is still small, but it surely will grow in time. To get help in English, there’s only one place #linux.